What is the line that passes through the points (4,7) and the slope of -0.75?
Answer by lwsshak3(11628)

What is the line that passes through the points (4,7) and the slope of -0.75?
**
Equation of a line: y=mx+b, m=slope, b=y-intercept
m=-0.75=-3/4
equation:
y=-3x/4+b
solve for b using coordinates of given point (4,7)
7=-3*4/4+b
b=10
equation of line: y=-3x/4+10